OpenStudy (anonymous):

Need Help Need Help Solve using elimination: 5x − 2y − 1= 0 2y − 3x = 5

OpenStudy (anonymous):

(-5, -5) (1, 4) (3, 7) (0, -1/2) its one of these

OpenStudy (anonymous):

Start by adding the equations: 5x-2y-1=0 +2y-3x=5 =5x-3x-1=5 simplify: 2x-1=5 solve for x 2x=6, x=3 now plug back into one of the original equations and solve for y.

OpenStudy (anonymous):

@Ihm ??

OpenStudy (anonymous):

So, you have two equations, one has -2y in it and the other has +2y in it. By adding the left sides and the right sides you can eliminate y. (5x − 2y − 1)+( 2y − 3x) = 0+5

OpenStudy (anonymous):

mhmmmm ok

OpenStudy (anonymous):

Simplify